OpenStudy (ray1998):

What are the zeroes of the function? Graph the function. y = x(x + 2)(x + 5) A. 0, 2, 5 B. -2, -5 C. -2, -5, 2 D. 0, -2, -5 Fan and medal will be given to whomever is most helpful :) thank you in advance!

Mehek (mehek14):

x + 2 = 0 x + 5 = 0 solve for x in both

random231 (random231):

yup mehek is right!

Mehek (mehek14):

and also x + 0 = 0

OpenStudy (ray1998):

Oh! Okay, well thank you :) you don't need the graphs? I was just about to post them? @Mehek14

Mehek (mehek14):

no lol

OpenStudy (ray1998):

Alright :) thank you very much guys!

Mehek (mehek14):

yw :)